Twenty metrics of liquidity, Solvency, and Profitability

The comparative financial statements of Automotive Solutions Inc. are as follows. The market price of Automotive Solutions Inc. common stock was $59 on December 31, 20Y8.

AUTOMOTIVE SOLUTIONS INC.
Comparative Income Statement
For the Years Ended December 31, 20Y8 and 20Y7
    20Y8     20Y7
Sales $6,137,475 $5,654,750
Cost of goods sold (2,124,300) (1,954,360)
Gross profit $4,013,175 $3,700,390
Selling expenses $(1,438,910) $(1,738,690)
Administrative expenses (1,225,745) (1,021,140)
Total operating expenses (2,664,655) (2,759,830)
Operating income $1,348,520 $940,560
Other revenue and expense:
    Other income 70,980 60,040
    Other expense (interest) (424,000) (233,600)
Income before income tax $995,500 $767,000
Income tax expense (119,500) (91,900)
Net income $876,000 $675,100


AUTOMOTIVE SOLUTIONS INC.
Comparative Statement of Stockholders’ Equity
For the Years Ended December 31, 20Y8 and 20Y7
20Y8 20Y7
Preferred
Stock
Common
Stock
Retained
Earnings
Preferred
Stock
Common
Stock
Retained
Earnings
Balances, Jan. 1 $800,000 $910,000 $3,891,725 $800,000 $910,000 $3,296,075
Net income 876,000 675,100
Dividends:
    Preferred stock (11,200) (11,200)
    Common stock (68,250) (68,250)
Balances, Dec. 31 $800,000 $910,000 $4,688,275 $800,000 $910,000 $3,891,725


AUTOMOTIVE SOLUTIONS INC.
Comparative Balance Sheet
December 31, 20Y8 and 20Y7
    Dec. 31, 20Y8     Dec. 31, 20Y7
Assets
Current assets:
Cash $1,363,760 $835,920
Temporary investments 2,064,060 1,385,240
Accounts receivable (net) 1,073,100 1,007,400
Inventories 803,000 613,200
Prepaid expenses 258,010 167,180
Total current assets $5,561,930 $4,008,940
Long-term investments 2,044,448 521,968
Property, plant, and equipment (net) 5,830,000 5,247,000
Total assets $13,436,378 $9,777,908
Liabilities
Current liabilities $1,738,103 $1,256,183
Long-term liabilities:
Mortgage note payable, 8%, due in 15 years $2,380,000 $0
Bonds payable, 8%, due in 20 years 2,920,000 2,920,000
Total long-term liabilities $5,300,000 $2,920,000
Total liabilities $7,038,103 $4,176,183
Stockholders' Equity
Preferred $0.70 stock, $50 par $800,000 $800,000
Common stock, $10 par 910,000 910,000
Retained earnings 4,688,275 3,891,725
Total stockholders' equity $6,398,275 $5,601,725
Total liabilities and stockholders' equity $13,436,378 $9,777,908

Instructions:

Determine the following measures for 20Y8. Round ratio values to one decimal place and dollar amounts to the nearest cent. For number of days' sales in receivables and number of days' sales in inventory, round intermediate calculations to the nearest whole dollar and final amounts to one decimal place. Assume there are 365 days in the year.

1. Working capital $
2. Current ratio
3. Quick ratio
4. Accounts receivable turnover
5. Days' sales in receivables days
6. Inventory turnover
7. Days' sales in inventory days
8. Debt ratio %
9. Ratio of liabilities to stockholders' equity
10. Ratio of fixed assets to long-term liabilities
11. Times interest earned times
12. Times preferred dividends earned times
13. Asset turnover
14. Return on total assets %
15. Return on stockholders’ equity %
16. Return on common stockholders’ equity %
17. Earnings per share on common stock $
18. Price-earnings ratio
19. Dividends per share of common stock $
20. Dividend yield %
Answer #1
1. Working capital = 5561930-1738103 = $ 38,23,827.00
2. Current ratio = 5561930/1738103 = 3.2
3. Quick ratio = (1363760+2064060+1073100)/1738103 = 2.6
4. Accounts receivable turnover = 6137475/((1073100+1007400)/2)) = 5.9
5. Days' sales in receivables (365/5.9) 61.9 days
6. Inventory turnover = 2124300/((803000+613200)/2)) = 3.0
7. Days' sales in inventory = 365/3 = 121.7 days
8. Debt ratio = 7038103/13436378 = 52.4% %
9. Ratio of liabilities to stockholders' equity = 7038103/6398275 = 1.1
10. Ratio of fixed assets to long-term liabilities = 5830000/7038103 = 0.8
11. Times interest earned = 1348520/424000 = 3.2 times
12. Times preferred dividends earned = 876000/11200 = 78.2 times
13. Asset turnover = 6137475/((13436378+9777908)/2) = 0.5
14. Return on total assets = 876000/((13436378+9777908)/2) = 7.5% %
15. Return on stockholders’ equity = 876000/((6398275+5601725))/2) = 3.7% %
16. Return on common stockholders’ equity = (876000-11200)/((910000+910000+4688275+3891725)/2)) = 16.6% %
17. Earnings per share on common stock = (876000-11200)/91000 = $              9.50
18. Price-earnings ratio = 59/9.5 = 6.2
19. Dividends per share of common stock = 68250/91000 = $              0.75
20. Dividend yield = .75/59 = 1.3% %